OpenStudy (anonymous):

Simplify: 9 1 __ + __ 14 4 ____________ 9 2 __ - ___ 7 21

Directrix (directrix):

Write the numerator 9/14 + 1/4 with a common denominator and add to a single fraction. Next, do the same with the denominator. Then, the division takes place.

Directrix (directrix):

What is a common denominator for these: 9/14 + 1/4

OpenStudy (anonymous):

28?

Directrix (directrix):

Yes.

Directrix (directrix):

9/14 = ?/28 1/4 = ?/28

OpenStudy (anonymous):

\[\frac{ 9 }{ 7 } - \frac{ 2 }{ 21 }\] \[\frac{ 9(3) }{ 7(3) } - \frac{ 2 }{ 21 }\] \[\frac{ 27 }{ 21 } - \frac{ 2 }{ 21 }\] \[\frac{ 25 }{ 21 }\]

Directrix (directrix):

I did not get 25/21. @xxxo What did you get?

OpenStudy (anonymous):

9/14 = 63/28 1/4= 2/28

OpenStudy (anonymous):

\[\frac{ 9 }{ 14 } + \frac{ 1 }{ 4 }\] \[\frac{ 9(2) }{ 14(2) } + \frac{ 1(7) }{ 4(7) }\] \[\frac{ 18 }{ 28 } + \frac{ 7 }{ 28 }\] \[\frac{ 25 }{ 28 }\]

Directrix (directrix):

9/14 = 63/28 1/4= 2/28 --------------- These are not correct.

Directrix (directrix):